OpenStudy (anonymous):

Solve the system 2x + 2y = -6 and 3x - 2y = 11 by using graph paper or graphing technology. What is the solution to the system? (-1, -7) (1, -4) (2, -1) (3, -2) i got d

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

plot each equation on the xy axis system I'd use something like desmos https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

the point where the two lines cross is the solution

OpenStudy (anonymous):

i got 1,-4

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

don't forget the parenthesis though

OpenStudy (triciaal):

to graph a line you need a minimum of 2 points the easiest to use what is x when y = 0 and what is y then x = 0 to find the solution as stated above find where the lines cross

OpenStudy (anonymous):

At the end of May, Janet told Sam that she has read 10 books this year and reads 2 books each month. Sam wants to catch up to Janet. He tracks his book reading with a table on his door. Using his table below, what month will Sam have read the same amount of books as Janet? Month Books June 3 July 6 August 9 -November -December -January -March i got november?

jimthompson5910 (jim_thompson5910):

This is throwing me off `Janet told Sam that she has read 10 books this year and reads 2 books each month` if she read 2 books each month, then she actually has read 2*12 = 24 books total for the year. How can they say she only read 10 books in a year?

OpenStudy (usukidoll):

I had a different approach.. that's why I was hesitant to answer it said at the end of May she told Same that she read 10 books this year and reads 2 books each month so it's at the beginning of the year... from January to May .. since the whole year hasn't be completed yet

jimthompson5910 (jim_thompson5910):

oh that could be a possibility

OpenStudy (triciaal):

@jim_thompson5910 will 2x-14 = 3x work where x is the month

OpenStudy (usukidoll):

I'm thinking that since it's at the end of May and we now know that it's 10 book and reads 2 a month January -2 Febuary 2 and so on until May

OpenStudy (usukidoll):

omg lack of sleep. It's probably starting at January and then at the end of May it was revealed that she read 10 books so far this year... She reads 2 a month so she was reading 2 books for 5 months.. January, Feb, March, April May

OpenStudy (usukidoll):

and he's like ok I'm going to catch up.... let me just start at June but instead of 2 a month, it's 3 a month

OpenStudy (usukidoll):

so he's like first month 3 books.. makes sense first month x 3 books 1 x 3 = 3

OpenStudy (usukidoll):

second month x 3 books 2 x 3 third month x 3 books 3 x 3 and then I'm assuming that he takes a break and restarts in November now fourth month x 3

OpenStudy (usukidoll):

So Sam's Chart was Month Books June 3 July 6 August 9 then we have November, so he didn't read for September and October that could be Sept 0 Oct 0 November (resumes reading) or something like that and I see the multiples of 3 being used too before his break

OpenStudy (triciaal):

-14 + 12 =-2 agree November

jimthompson5910 (jim_thompson5910):

@triciaal you're on the right path it should be 2x + 10 = 3x though x = number of months since May eg: x = 0 represents may x = 1 represents june x = 2 represents july etc

OpenStudy (usukidoll):

I also got November, but in a different style .

OpenStudy (usukidoll):

I don't like these problems too much.. worse than Mathematical Induction

OpenStudy (triciaal):

@jim_thompson5910 thanks yes that works 2

jimthompson5910 (jim_thompson5910):

The answer isn't Nov

OpenStudy (usukidoll):

triciaal also got November.. so why is it not November?

jimthompson5910 (jim_thompson5910):

oh wait the equation isn't 2x + 10 = 3x solve for x to get 2x + 10 = 3x 2x + 10-2x = 3x-2x 10 = x x = 10 may is the 5th month 10 months from there is 5+10 = 15 the 15th month is really the 3rd month since 15-12 = 3

jimthompson5910 (jim_thompson5910):

the table confirms it

OpenStudy (triciaal):

so it would be March?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (triciaal):

the table confirmed

OpenStudy (triciaal):

@UsukiDoll I had it wrong because I did 12 months not from May 5- 2 = 3 the 3rd month March

OpenStudy (usukidoll):

@triciaal I figured out my mistake . Apparently Sam's Chart blended into the Multiple choices so I assumed that Sam took breaks when in fact he didn't so he did started in June by reading 3 books a month until Sam can catch up with her. So Sam was the multiple of 3 and she was the multiple of 2. With enough reading, both of them will catch up with each other in March for reading 30 books.

OpenStudy (usukidoll):

Janet was a multiple of 2 and started reading in January.. so by March she will have 30 2 4 6 8 10 12 14 16 18 20 22 24 26 28 30 (1 year and 3 months_ Sam was a multiple of 3 and started in June 0 0 0 0 0 3 6 9 12 15 18 21 24 27 30 (1 year and 3 months) Although he had a late start, adding one more book in his reading list proves to be successful later on.